OpenStudy (anonymous):

solve the equation. 3(x-2) +5 =x+2(x+1)

OpenStudy (aroub):

3x-6+5=x+2x+2 3x-1=3x+2 3x-3x=1+2 0=3 => False or no solution

OpenStudy (anonymous):

Thank you.